OpenStudy (anonymous):

What is OpenStudy Feedback for? For example Mathematics is for studying math question.

OpenStudy (anonymous):

For "Feedback"

OpenStudy (aravindg):

You provide your ideas and suggestions here-as the name suggests.

mathslover (mathslover):

\(\LARGE{\underline{\color{blue}{\textbf{What is}} \space \color{Orange}{\mathbb{OpenStudy}} \space \color{Red}{\mathsf{Feedback} }\space \textbf{for?} }}\) OpenStudy Feedback is a group which basically allows user to \(\color{blue}{ \underline{ \textbf{give feedback/suggestions}}}\) about the : \(\color{blue}{\bullet}\) Functions/Features of OpenStudy \(\color{blue}{\bullet}\) Help being given here \(\color{blue}{\bullet}\)Bugs while working on OpenStudy It also gives the users \(\color{red}{ \underline{ \textbf{a chance to :}}}\) \(\color{red}{\bullet}\) Wish their friends of OpenStudy on special occasions like Birthday etc. \(\color{red}{\bullet}\)Post their latest achievements and milestones so that others get a chance to get inspiration from them and also to acknowledge them on their success. \(\color{red}{\bullet}\)Post their interesting stories/experience while working on OpenStudy. It is also a place for the administrators/moderators of OpenStudy to \(\color{green}{ \underline{ \textbf{make announcements related to :}}}\) \(\color{green}{\bullet}\) New features implemented. \(\color{green}{\bullet}\) Changes in the Code-of-Conduct / Privacy Policy / Terms and Conditions. \(\color{green}{\bullet}\) Announcing new mods/ambis . \(\color{green}{\bullet}\) Announcement of new programs like Ambis new program. \(\color{green}{\bullet}\) Future downtimes or past ones. \(\bf \color{blue}{----------- \times ---------------\times --------}\)